PLEASE HELP WILL UPVOTE

The company used a job order cost system to accumulate product costs. Use the T-Charts below to help answer the questions below.

RAW MATERIALS INVENTORY

65,450

451,500

440,900

BAL: 76,050

WORK IN PROCESS INVENTORY

34,960

440,900

411,600

172,020

217,140

BAL: 453,420

FINISHED GOODS INVENTORY

311,500

411,600

425,740

BAL: 297,360

  1. Raw Materials purchased during 2018?
  2. Direct Materials Entered into production during March?
  3. Cost of Goods Manufactured during March?
  4. Cost of Goods Sold during March? * (REQUIRES EXPLANATION)
  5. Conversion Costs Added into the factory during March? * (REQUIRES EXPLANATION)
  6. Production Costs reported on Balance sheet ay March 31, 2018? *(REQUIRES EXPLANATION)
Answer #1

a) Raw material purchase = $451500

b) Direct material entered into production = 440900

c) Cost of goods manufactured = 411600

d) Cost of goods sold = Beginning finished goods+Cost of goods manufactured-Ending finished goods

= 311500+411600-297360

Cost of goods sold = 425740

e) Conversion Cost = Direct labor+Manufacturing overhead = 172020+217140 = $389160

f) Production costs reported on balance sheet

Raw material 76050
Work in process 453420
Finished goods 297360
Total 826830